# Fano threefolds in positive characteristic III

### v3

@@ 5.1. The Picard group and the canonical divisor
Thus we obtain $c_{2}(X)\cdot H_{i}=\frac{24}{r_{i}}$, which completes the proof for the case when $R_{i}$ is of type $E_{2}$.
Assume that $R_{i}$ is of type $E_{3}$ or $E_{4}$. It follows from [TanII, Proposition 3.22] that $\mu_{i}=1$, $(-K_{X})^{2}\cdot D_{i}=2$, $D_{i}^{3}=2$, $D_{i}$ is isomorphic to a possibly singular quadric surface in $\mathbb{P}^{3}$, $\mathcal{O}_{X}(D_{i})|_{D_{i}}\simeq\mathcal{O}_{\mathbb{P}^{3}}(-1)|_{D_{i}}$, and $\ell_{i}$ is a line on $\mathbb{P}^{3}$ contained in $D_{i}$ (Theorem 4.8). In particular, $D_{i}\cdot\ell_{i}=-1$. We have $-K_{X}+D_{i}\sim-f_{i}^{*}K_{Y_{i}}\sim f_{i}^{*}(r_{i}L_{i})=r_{i}H_{i}$. We then get
$c_{2}(X)\cdot H_{i}=\frac{1}{r_{i}}(24+c_{2}(X)\cdot D_{i}).$
Let us compute $c_{2}(X)\cdot D_{i}$. By Lemma 3.8(4), we have
It holds that $\chi(D_{i},\mathcal{O}_{D_{i}})=h^{0}(D_{i},\mathcal{O}_{D_{i}})=1$ and $\chi(D_{i},\mathcal{O}_{D_{i}}(D_{i}))=\chi(D_{i},\mathcal{O}_{\mathbb{P}^{3}}(-1)|_{D_{i}})=0$. Therefore, we obtain
$c_{2}(X)\cdot D_{i}=6+0-4-2=0,$
which implies $c_{2}(X)\cdot H_{i}=\frac{24}{r_{i}}$. This completes the proof for the case when $R_{i}$ is of type $E_{3}$ or $E_{4}$.
Assume that $R_{i}$ is of type $E_{5}$. It follows from [TanII, Proposition 3.22] that $\mu_{i}=1$, $D_{i}\simeq\mathbb{P}^{2}$, $D_{i}^{3}=4$, $(-K_{X})^{2}\cdot D_{i}=1$, $\mathcal{O}_{X}(D_{i})|_{X_{i}}\simeq\mathcal{O}_{\mathbb{P}^{2}}(-2)$, and $\ell_{i}$ is a line on $D_{i}$. Hence
$(-2K_{X}+D_{i})\cdot\ell_{i}=-2K_{X}\cdot\ell_{i}+D_{i}\cdot\ell_{i}=2-2=0,$
which implies $-2K_{X}+D_{i}\sim r_{i}H_{i}$. As before, we can write
$c_{2}(X)\cdot H_{i}=\frac{1}{r_{i}}(2\cdot 24+c_{2}(X)\cdot D_{i}).$
By Lemma 3.8(4), we have
It holds that $\chi(D_{i},\mathcal{O}_{D_{i}})=1$ and $\chi(D_{i},\mathcal{O}_{X}(D_{i})|_{D_{i}})=0$. We obtain
$c_{2}(X)\cdot D_{i}=6+0-8-1=-3,$
and hence $c_{2}(X)\cdot H_{i}=\frac{45}{r_{i}}$. Thus (3) holds. ∎
@@
and $\deg\Delta_{f_{i}}>0$. In particular, $(c_{2}(X)\cdot H_{i})\geq 7$. Thus (1) holds.
Let us show (2). If $R_{i}$ not of type $D_{1}$, then these assertions follow from (1) and Lemma 5.4. Assume that $R_{i}$ is of type $D_{1}$. Then Lemma 5.4 implies
@@
$a:=|{\operatorname{Pic}}(X)/(\mathbb{Z}H_{1}\oplus\mathbb{Z}H_{2})|,$
which is the cardinality of the set ${\operatorname{Pic}}(X)/(\mathbb{Z}H_{1}\oplus\mathbb{Z}H_{2})$. Then the following hold.
@@
    $H_{1}$ and $H_{2}$ are linearly equivalent over $\mathbb{R}$ in $({\operatorname{Pic}}\,X)\otimes_{\mathbb{Z}}\mathbb{R}/\equiv$. Moreover, $a\in\mathbb{Z}_{>0}$.
@@
for $\lambda_{1},\lambda_{2}\in\mathbb{R}$. By $H_{1}=f^{*}_{1}L_{1}$, we have $H_{1}\cdot\ell_{1}=0$. Hence we obtain $\lambda_{2}H_{2}\cdot\ell_{1}=0$. Since $f_{2}(\ell_{1})$ is a curve, we have $H_{2}\cdot\ell_{1}>0$. We then get $\lambda_{2}=0$, which implies $\lambda_{1}=0$. Therefore, $H_{1}$ and $H_{2}$ are linearly independent over $\mathbb{R}$. Since $\mathbb{Z}H_{1}\oplus\mathbb{Z}H_{2}$ is a free subgroup of rank $2$ of ${\operatorname{Pic}}\,X\simeq\mathbb{Z}^{\oplus 2}$, the quotient group
@@
Thus we obtain $H-bH_{2}\in\mathrm{Ker}(-\cdot\ell_{1})=\mathrm{Im}(f^{*}_{1})=\mathbb{Z}H_{1}$ by Theorem 4.11. Then we can write $H\sim cH_{1}+bH_{2}$ for some $c\in\mathbb{Z}$. Thus (2) and (3) holds.
@@
Set $R_{1}:=R$ and $a:=|{\operatorname{Pic}}(X)/(\mathbb{Z}H_{1}\oplus\mathbb{Z}H_{2})|\in\mathbb{Z}_{>0}$ (Lemma 5.6(1)). We use Notation 5.1. By $-K_{Y}=f_{*}(-K_{X})$, $-K_{Y}$ is big. This, together with $\rho(Y)=1$, implies that $-K_{Y}$ is ample. Suppose $r_{Y}=1$. It is enough to derive a contradiction. Recall that $|-K_{Y}|$ is not very ample [TanII, Theorem 1.1]. By [TanI, Theorem 1.1], we obtain $(-K_{Y})^{3}\in\{2,4\}$. It follows from $(-K_{X})^{3}\in 2\mathbb{Z}$ and $0<(-K_{X})^{3}<(-K_{Y})^{3}$ (Lemma 3.13) that $(-K_{Y})^{3}=4$ and $(-K_{X})^{3}=2$. Let $\varphi_{|-K_{X}|}:X\dashrightarrow\mathbb{P}^{h^{0}(X,-K_{X})-1}$ be the rational map induced by the complete linear system $|-K_{X}|$. If $\dim({\operatorname{Im}}\,\varphi_{|-K_{X}|})\geq 2$, then our condition $\rho(X)=2$ contradicts [TanI, Proposition 6.8]. Hence we may assume that $\dim({\operatorname{Im}}\,\varphi_{|-K_{X}|})=1$. In this case, the mobile part $M$ of $|-K_{X}|$ is base point free and the induced morphism $\varphi_{|M|}:X\to\mathbb{P}^{1}$ is the contraction of the extremal ray $R_{2}$ of type D with $d_{2}:=(-K_{X})^{2}\cdot H_{2}=1$ [TanI, Proposition 4.1]. In particular, $\mu_{2}=1$. Then Lemma 5.4 and Lemma 5.6 imply
@@
Since $X$ is imprimitive, $X$ has an extremal ray of type $E_{1}$. In what follows, we assume that $R_{1}$ is of type $E_{1}$. Set $a:=|{\operatorname{Pic}}(X)/(\mathbb{Z}H_{1}\oplus\mathbb{Z}H_{2})|\in\mathbb{Z}_{>0}$ (Lemma 5.6(1)). It is enough to show $a=1$ (cf. Step 1 of the primitive case).
@@
    $H_{1}\cdot c_{2}(X)\leq\begin{cases}25\qquad(\text{if }a\leq 3)\\ 23\qquad(\text{if }\text{if }a\leq 2)\\ 17\qquad(\text{if }a=1).\end{cases}$
@@
In what follows, we shall separate the argument depending on the value of $r_{1}$. Note that $r_{1}\in\{2,3,4\}$ (Proposition 5.8).
Assume $r_{1}=4$. In this case, we have $L_{1}^{3}=1$. This, together with (5.10.1), implies
@@
Assume $r_{1}=3$. In this case, we have $L_{1}^{3}=2$. This, together with (5.10.1), implies
@@
Assume $r_{1}=2$. In this case, we have $1\leq L_{1}^{3}\leq 5$. This, together with (5.10.1), implies
@@
This is absurd. Thus $R_{2}$ is not of type $E_{2}$.
@@
If $r_{1}=4$, then $L_{1}^{3}=1$, and hence $2231\leq 6+(4-\frac{3}{a})^{2}<6+16=$ 22, which is absurd. If $r_{1}=3$, then $L_{1}^{3}=2$, and hence $2631\leq 8+(3-\frac{3}{a})^{2}\cdot 2<8+9\cdot 2=$ 26, which is impossible. If $r_{1}=2$, then $1\leq L_{1}^{3}\leq 5$, and hence $831\leq 12+(2-\frac{3}{4})^{2}\cdot L_{1}^{3}\leq 12+\frac{25}{16}\cdot 5<12+$ 8, which is a contradiction. We then obtain $a\neq 4$.
@@ 5.2. Case $C-C$
We use Notation 5.1. Assume that $R_{1}$ is of type $C$ and $R_{2}$ is of type $C$. In particular, for each $i\in\{1,2\}$, $f_{i}:X\to Y_{i}=\mathbb{P}^{2}$ is of type $C$. Set
@@
Let $\widetilde{f}:X\to f(X)$ be the induced morphism, where $f(X)$ denotes the scheme-theoretic image. For the $i$-th projection $\pi_{i}:\mathbb{P}^{2}\times\mathbb{P}^{2}\to\mathbb{P}^{2}$, set $M_{i}:=\pi^{*}L_{i}$. Note that $\{M_{1},M_{2}\}$ is a free $\mathbb{Z}$-linear basis of ${\operatorname{Pic}}(\mathbb{P}^{2}\times\mathbb{P}^{2})$.
@@ 5.2.2. Case $\deg(\widetilde{f})=2$
Applying the elementary operations to the $3\times 3$-matrix $(a_{ij})$, we may assume that $W$ is one of $W_{1},W_{2}$, and $W_{3}$ (note that we can apply row and column elementray operations).
@@
We then get $deg(\tilde{f})=2$ and $\mu_{1}=\mu_{2}=1$. Hence (1) holds. By Lemma 5.14, $(-K_{X})^{3}=6(\mu_{1}^{2}+\mu_{2}^{2})=12.$ Thus (2) holds. Furthermore, we have
@@
Then the assertion (3) follows from Lemma 5.16(2).
@@
    $R_{1}$ is of type $C_{1}$, $R_{2}$ is of type $C_{1}$, and one of the following holds.
@@
    $R_{1}$ is of type $C_{1}$, $R_{2}$ is of type $C_{2}$, $(-K_{X})^{3}=30$, and $X$ is isomorphic to a hypersurface of $\mathbb{P}^{2}\times\mathbb{P}^{2}$ of bidegree $(1,2)$.
@@
    $R_{1}$ is of type $C_{2}$, $R_{2}$ is of type $C_{2}$, $(-K_{X})^{3}=48$, and $X$ is isomorphic to $W$.
@@
If one of the contractions is not generically smooth, then (2) holds by Proposition 4.17(5). Hence we may assume that each contraction is generically smooth. Then the assertion follows from Lemma 5.14, Lemma 5.15, and Lemma 5.17. ∎

### v4

@@ 5.1. The Picard group and the canonical divisor
Thus we obtain
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}c_{2}(X)\cdot H_{i}=\frac{\epsilon}{r_{i}}(24+ac_{2}(X)\cdot D_{i})=\frac{1}{r_{i}}(24+0)=\frac{24}{r_{i}},}$
which completes the proof for the case when $R_{i}$ is of type $E_{2}$.
Assume that $R_{i}$ is of type $E_{3}$ or $E_{4}$. It follows from [TanII, Proposition 3.22] that $\mu_{i}=1$, $(-K_{X})^{2}\cdot D_{i}=2$, $D_{i}^{3}=2$, $D_{i}$ is isomorphic to a possibly singular quadric surface in $\mathbb{P}^{3}$, and $\mathcal{O}_{X}(D_{i})|_{D_{i}}\simeq\mathcal{O}_{\mathbb{P}^{3}}(-1)|_{D_{i}}$. It holds that $\chi(D_{i},\mathcal{O}_{D_{i}})=h^{0}(D_{i},\mathcal{O}_{D_{i}})=1$ and $\chi(D_{i},\mathcal{O}_{D_{i}}(D_{i}))=\chi(D_{i},\mathcal{O}_{\mathbb{P}^{3}}(-1)|_{D_{i}})=0$. Therefore, we obtain
$=6+0-4-2=0.$
We then get
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}c_{2}(X)\cdot H_{i}=\frac{\epsilon}{r_{i}}(24+ac_{2}(X)\cdot D_{i})=\frac{1}{r_{i}}(24+0)=\frac{24}{r_{i}},}$
which completes the proof for the case when $R_{i}$ is of type $E_{3}$ or $E_{4}$.
Assume that $R_{i}$ is of type $E_{5}$. It follows from [TanII, Proposition 3.22] that $\mu_{i}=1$, $D_{i}\simeq\mathbb{P}^{2}$, $D_{i}^{3}=4$, $(-K_{X})^{2}\cdot D_{i}=1$, and $\mathcal{O}_{X}(D_{i})|_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}D_{i}}}\simeq\mathcal{O}_{\mathbb{P}^{2}}(-2)$. It holds that $\chi(D_{i},\mathcal{O}_{D_{i}})=1$ and $\chi(D_{i},\mathcal{O}_{X}(D_{i})|_{D_{i}})=0$. We obtain
$=6+0-8-1=-3.$
We then get
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}c_{2}(X)\cdot H_{i}=\frac{\epsilon}{r_{i}}(24+ac_{2}(X)\cdot D_{i})=\frac{2}{r_{i}}(24+\frac{1}{2}\cdot(-3))=\frac{45}{r_{i}},}$
which completes the proof for the case when $R_{i}$ is of type $E_{5}$. Thus (3) holds. ∎
@@
and $\deg\Delta_{f_{i}}>0$. In particular, $c_{2}(X)\cdot H_{i}\geq 7$. Thus (1) holds.
Let us show (2). If $R_{i}$ is not of type $D_{1}$, then these assertions follow from (1) and Lemma 5.4. Assume that $R_{i}$ is of type $D_{1}$. Then Lemma 5.4(2) implies
@@
$a:=|{\operatorname{Pic}}(X)/(\mathbb{Z}H_{1}+\mathbb{Z}H_{2})|,$
which is the cardinality of the set ${\operatorname{Pic}}(X)/(\mathbb{Z}H_{1}+\mathbb{Z}H_{2})$. Then the following hold.
@@
    $H_{1}$ and $H_{2}$ are linearly independent over $\mathbb{R}$ in $({\operatorname{Pic}}\,X)\otimes_{\mathbb{Z}}\mathbb{R}/\equiv$. Moreover, $a\in\mathbb{Z}_{>0}$ and $\mathbb{Z}H_{1}\oplus\mathbb{Z}H_{2}=\mathbb{Z}H_{1}+\mathbb{Z}H_{2}$.
@@
for $\lambda_{1},\lambda_{2}\in\mathbb{R}$. By $H_{1}=f^{*}_{1}L_{1}$, we have $H_{1}\cdot\ell_{1}=0$. Hence we obtain $\lambda_{2}H_{2}\cdot\ell_{1}=0$. Since $f_{2}(\ell_{1})$ is a curve, we have $H_{2}\cdot\ell_{1}>0$. We then get $\lambda_{2}=0$, which implies $\lambda_{1}=0$. Therefore, $H_{1}$ and $H_{2}$ are linearly independent over $\mathbb{R}$. In particular, we get $\mathbb{Z}H_{1}\oplus\mathbb{Z}H_{2}=\mathbb{Z}H_{1}+\mathbb{Z}H_{2}$. Since $\mathbb{Z}H_{1}\oplus\mathbb{Z}H_{2}$ is a free subgroup of rank $2$ of ${\operatorname{Pic}}\,X\simeq\mathbb{Z}^{\oplus 2}$, the quotient group
@@
Thus we obtain $H-bH_{2}\in\mathrm{Ker}(-\cdot\ell_{1})=\mathrm{Im}(f^{*}_{1})=\mathbb{Z}H_{1}$ by Theorem 4.11. Then we can write $H\sim cH_{1}+bH_{2}$ for some $c\in\mathbb{Z}$. Thus (2) and (3) hold.
@@
Set $R_{1}:=R$ and $a:=|{\operatorname{Pic}}(X)/(\mathbb{Z}H_{1}\oplus\mathbb{Z}H_{2})|\in\mathbb{Z}_{>0}$ (Lemma 5.6(1)). We use Notation 5.1. By $-K_{Y}=f_{*}(-K_{X})$, $-K_{Y}$ is big. This, together with $\rho(Y)=1$, implies that $-K_{Y}$ is ample. Suppose ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}r_{1}}=r_{Y}=1$. It is enough to derive a contradiction. Recall that $|-K_{Y}|$ is not very ample [TanII, Theorem 1.1(3)]. By [TanI, Theorem 1.1], we obtain $(-K_{Y})^{3}\in\{2,4\}$. We have $(-K_{X})^{3}\in 2\mathbb{Z}$ by the Riemann-Roch theorem $\chi(X,-K_{X})=\frac{(-K_{X})^{3}}{2}+3$ [TanI, Corollary 2.6]. These, together with $0<(-K_{X})^{3}<(-K_{Y})^{3}$ (Lemma 3.13), imply that $(-K_{Y})^{3}=4$ and $(-K_{X})^{3}=2$. Let $\varphi_{|-K_{X}|}:X\dashrightarrow\mathbb{P}^{h^{0}(X,-K_{X})-1}$ be the rational map induced by the complete linear system $|-K_{X}|$. If $\dim({\operatorname{Im}}\,\varphi_{|-K_{X}|})\geq 2$, then our condition $\rho(X)=2$ contradicts [TanI, Proposition 6.8]. Hence we may assume that $\dim({\operatorname{Im}}\,\varphi_{|-K_{X}|})=1$. In this case, the mobile part $M$ of $|-K_{X}|$ is base point free and the induced morphism $\varphi_{|M|}:X\to\mathbb{P}^{1}$ is the contraction of the extremal ray $R_{2}$ of type D with $d_{2}:=(-K_{X})^{2}\cdot H_{2}=1$ [TanI, Proposition 4.1] (note that the equality $-K_{X}^{2}\cdot E=1$ claimed in [TanI, Proposition 4.1(3)] is a typo; the correct equality is $(-K_{X})^{2}\cdot E=1$, as given in the last line of its proof). In particular, $\mu_{2}=1$. Then Lemma 5.4 and Lemma 5.6 imply
@@
Since $X$ is imprimitive, $X$ has an extremal ray of type $E_{1}$. In what follows, we assume that $R_{1}$ is of type $E_{1}$. Set $a:=|{\operatorname{Pic}}(X)/(\mathbb{Z}H_{1}\oplus\mathbb{Z}H_{2})|\in\mathbb{Z}_{>0}$ (Lemma 5.6(1)). It is enough to show $a=1$ (cf. Step 1 of the primitive case). Recall that $2\leq r_{1}\leq 4$ (Proposition 5.8).
@@
    $H_{1}\cdot c_{2}(X)\leq\begin{cases}25\qquad(\text{if }a\leq 3)\\ 23\qquad(\text{if }a\leq 2)\\ 17\qquad(\text{if }a=1).\end{cases}$
@@
In what follows, we shall separate the argument depending on the value of $r_{1}$.
Assume $r_{1}=4$. In this case, we have $Y_{1}\simeq\mathbb{P}^{3}$ and $L_{1}^{3}=1$ [TanI, Theorem 2.18]. This, together with (5.10.1), implies
@@
Assume $r_{1}=3$. In this case, $Y_{1}$ is a smooth quadric hypersurace in $\mathbb{P}^{4}$ and we have $L_{1}^{3}=2$ [TanI, Theorem 2.18]. This, together with (5.10.1), implies
@@
Assume $r_{1}=2$. In this case, we have $1\leq L_{1}^{3}\leq 5$ [TanII, Remark 4.9]. This, together with (5.10.1), implies
@@
This is absurd. Thus $R_{2}$ is not of type ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}E_{1}}$.
@@
If $r_{1}=4$, then $L_{1}^{3}=1$, and hence $2231\leq 6+(4-\frac{3}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}4}})^{2}<6+16=$ 22, which is absurd. If $r_{1}=3$, then $L_{1}^{3}=2$, and hence $2631\leq 8+(3-\frac{3}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}4}})^{2}\cdot 2<8+9\cdot 2=$ 26, which is impossible. If $r_{1}=2$, then $1\leq L_{1}^{3}\leq 5$, and hence $831\leq 12+(2-\frac{3}{4})^{2}\cdot L_{1}^{3}\leq 12+\frac{25}{16}\cdot 5<12+$ 8, which is a contradiction. We then obtain $a\neq 4$.
@@ 5.2. Case $C-C$
We use Notation 5.1. Assume that $R_{1}$ is of type $C$ and $R_{2}$ is of type $C$. In particular, for each $i\in\{1,2\}$, $f_{i}:X\to Y_{i}=\mathbb{P}^{2}$ is of type $C$ (Theorem 4.17). Set
@@
Let $\widetilde{f}:X\to f(X)$ be the induced morphism, where $f(X)$ denotes the scheme-theoretic image. For the $i$-th projection $\pi_{i}:\mathbb{P}^{2}\times\mathbb{P}^{2}\to\mathbb{P}^{2}$, set $M_{i}:=\pi_{i}^{*}L_{i}$. Note that $\{M_{1},M_{2}\}$ is a free $\mathbb{Z}$-linear basis of ${\operatorname{Pic}}(\mathbb{P}^{2}\times\mathbb{P}^{2})$.
@@ 5.2.2. Case $\deg(\widetilde{f})=2$
Applying the elementary operations to the $3\times 3$-matrix $(a_{ij})$, we may assume that $W$ is one of $W_{1},W_{2}$, and $W_{3}$ (note that we can apply row and column elementary operations). Thus (1) holds.
@@
We then get $\deg(\tilde{f})=2$ and $\mu_{1}=\mu_{2}=1$. Hence (1) holds. By Lemma 5.14, $(-K_{X})^{3}=6(\mu_{1}^{2}+\mu_{2}^{2})=12.$ Thus (2) holds. Furthermore, we have
@@
Then the assertion (3) follows from Lemma 5.16(2), because $f_{1}$ commutes with the projection ${\rm pr}_{1}$, i.e., $f_{1}:X\xrightarrow{\widetilde{f}}f(X)=W\hookrightarrow\mathbb{P}^{2}\times\mathbb{P}^{2}\xrightarrow{{\rm pr}_{1}}\mathbb{P}^{2}$ (Notation 5.12).
@@
    $R_{1}$ is of type $C_{1}$, $R_{2}$ is of type $C_{1}$, and one of the following holds (No. 2-6).
@@
    $R_{1}$ is of type $C_{1}$, $R_{2}$ is of type $C_{2}$, $(-K_{X})^{3}=30$, and $X$ is isomorphic to a hypersurface of $\mathbb{P}^{2}\times\mathbb{P}^{2}$ of bidegree $(1,2)$ (No. 2-24).
@@
    $R_{1}$ is of type $C_{2}$, $R_{2}$ is of type $C_{2}$, $(-K_{X})^{3}=48$, and $X$ is isomorphic to $W$ (No. 2-32).
@@
If one of the contractions is not generically smooth, then (2) holds by Theorem 4.2(2). Hence we may assume that each contraction is generically smooth. Then the assertion follows from Lemma 5.14, Lemma 5.15, and Lemma 5.17. ∎